# Interchangeable Token Embeddings for Extendable Vocabulary and Alpha-Equivalence

### v2

@@ 5 Experiments
Experimental setup
We use a transformer encoder-decoder architecture in all experiments. We always use the same embedding size in both encoder and decoder due to weight tying. We use the RoPE (Su et al., 2023) as the positional encoding method in both encoder and decoder unless otherwise noted. The hyperparameter settings are given in Table 4 in Section A.3.
Baselines
We train three types of baseline models with traditional embeddings: the first one on the original dataset, the second one on a dataset with the same parameters but using a larger vocabulary size, and the third one on the original dataset but using a data augmentation strategy. Specifically, for the third baseline, the number of interchangeable token embeddings matches that of the test set, and we apply random alpha-renaming at each forward pass during training. This ensures that the model is exposed to all tokens in the test set, but the number of unique interchangeable tokens the model sees in each sample remains limited as in the training set. Note that this is an internal baseline that doesn’t exist in the literature to the best of our knowledge.
@@ 5.1 Copying with Extendable Vocabulary
Figure 2: Two annotated heatmaps visualizing the test-set edit distance between prediction and ground truth in copying task with extendable vocabulary. Both heatmaps share the same y-axis. The green box represents the number of unique characters (y-axis) and the maximum length (x-axis) in the training dataset. The lower triangular part of each heatmap, shown in gray hatch pattern, represents the impossible combinations of length and unique character count. Each point represents the average error over test samples with a particular sequence length and unique number of character tokens. The baseline approach (on the left), using ubiquitously utilized fixed (learned) token embeddings, cannot extrapolate to vocabulary expansions. The proposed method (on the right) enables generalization to larger vocabulary sizes at longer sequence lengths, compared to what is observed during training.
We introduce a new toy problem designed to evaluate the vocabulary generalization capabilities of our embedding method. We create various training datasets that contain 10 million random strings with a limited vocabulary size. A string is given as input, and the model is expected to produce the input string exactly via autoregressive generation. This embodies a helpful toy problem for our method because all tokens are interchangeable, barring the special tokens (start/end). In these experiments, we expect the model to generalize to larger vocabulary sizes unseen during training. We generate the predictions using greedy sampling in this subsection.
Evaluation method
We use the edit distance between the prediction and the ground truth as our evaluation metric. To generate the evaluation datasets (validation and test splits), we create 100 samples for each possible combination of unique character count and string length, starting from a minimum of 3. Consequently, the total evaluation dataset is arranged in a matrix in which the rows represent unique character count in the string and the columns represent the string length. This matrix is upper triangular since the unique character count cannot exceed the string length. For random embeddings, we repeat the evaluation 10 times and report the average. To evaluate up to the string length of 30 in this setup, $10\times 100\times 406=406000$ predictions are required, where $406$ is the number of upper triangular elements in a $28\times 28$ matrix. To minimize the impact of random factors, we train each model three times and report the results only for the best.
5.1.1 Generalization to larger vocabularies
5.1.2 Generalization to larger vocabularies and lengths
We create a dataset consisting of 10 million strings whose lengths vary between 5 and 10 with at most 5 unique characters. We evaluate on the same validation set as before, expecting the model to generalize to both longer lengths and larger vocabulary sizes. In Appendix A.3.1, we perform a hyperparameter search over random embedding methods, $d_{\beta}$ values, and whether $f_{bn}$, $f_{fn}$, AdaCos are enabled.
5.1.3 Sensitivity to randomness in embeddings
To reduce the computational cost of evaluation in the next experiments, we generate 10 random embeddings, sort them by their cross entropy loss on the evaluated dataset, and use the median one. We find that this serves as a decent proxy for the average performance. Across the validation set evaluations of all 277 models, the percent difference in edit distance between this median method and the real mean is 1.4% on average (meaning that the result from the median method is worse), and 9.1% if we consider the absolute differences.
5.1.4 Scaling up
@@ 5.2 LTL Solving
In this section, we train models on the LTLRandom35 dataset from DeepLTL (Hahn et al., 2021) and other synthetic datasets created with the same method. To evaluate the correctness of the generated formulae, we utilize spot framework version 2.11.6 (Duret-Lutz et al., 2022). We use tree-positional encoding (Shiv and Quirk, 2019) in the encoder and RoPE (Su et al., 2023) in the decoder. We generate predictions using beam search with a beam size of 3 in this section.
Baselines
We trained all of the baseline models from scratch. For the first type of baseline, we aimed to reproduce the results from Hahn et al. (2021). Hence, we used the best hyperparameters they reported (Appendix A.3). Unlike Hahn et al. (2021), we experimented with RoPE (in the decoder) and AdaCos, but did not observe a noteworthy improvement on the validation set.  [^1] Using RoPE in the decoder increased the ratio of correct predictions from 97.8% to 98.0% on the validation set. Introducing AdaCos in addition to RoPE increased this value to 98.2%. After determining the best baseline model on the validation set, we evaluated it on the test split of LTLRandom35 and obtained a correct rate of $98.2\%$ against the $98.5\%$ reported by Hahn et al. (2021).
5.2.1 Dataset Perturbations
Table 2: Evaluation of the baselines and our method trained on different versions of LTLRandom35. The alpha-renaming baseline was trained using 5 AP embeddings since vocabulary generalization is not valuated here. First two columns denote the training dataset and the model. Next two columns indicate the ratio of the correct predictions and exact matches on 99,989 test set samples as evaluated by spot. Last three columns display mean alpha-covariance values for varying atomic proposition (AP) counts, evaluated on all alpha-equivalent variants of 1000 test samples.
@@ 5.2.1 Dataset Perturbations
While the original model performs significantly worse under perturbation, both alpha-renaming and proposed models match the baseline performance in correctness ratio despite perturbation. This observation suggests that these modifications introduce a robust inductive bias that makes the models resistant to perturbations in the data. A minor decrease in the ratio of exact matches is noted, but this may signify less overfitting and a better bias-variance tradeoff in the larger context. Appendix A.4 continues this experiment with limited amount of training samples instead of perturbations.
@@ 5.2.2 Alpha-Covariance
Intuitively, when alpha-covariance is $1$, the model is unaffected by all alpha-conversions in ${\mathbb{P}}$. An alpha-covariance of 00 indicates that $|{\mathbb{U}}|=|{\mathbb{P}}|$, i.e., the model’s prediction for each alpha-equivalent pair is unique after undoing the alpha-conversion. This is unwanted because alpha-conversions should not change the semantic meaning. Thanks to the embedding randomization in our method, an alpha-conversion does not necessarily change the embeddings, and conversely, there are multiple ways to embed the same input due to randomness.
For the proposed method, we generate the random embeddings once at the start of an evaluation run using the heuristic explained in Section 5.1.3. Thus, alpha-conversions in this experiment are equivalent to shuffling the random embeddings in our method. As a result, the alpha-covariance measures our model’s robustness against differences in random embeddings.
@@ 5.2.3 Generalization
The test dataset for this experiment contains at most 100 formula-trace pairs for each combination of AP count and formula length, whose maximum is 50 instead of 35. We report the results for our model (using Hypercube Vertices, $d_{\beta}=5$) and the three baselines in Figure 4. The first baseline uses the same training dataset, whereas the second baseline uses a new LTL dataset with 10 APs, which we create using the same method as LTLRandom35. For the third baseline, we train a fixed embedding model with 10 APs using the same 5 AP dataset but we shuffle the AP embeddings in each forward pass during training.
Figure 4: Heatmap visualizing the ratio of correct predictions on a special test set. The brightness of the color depends on the sample size, with full brightness representing 100 samples. The dashed white box represents the boundaries of the training dataset.
Discussion
@@
Despite seeing only 5 APs during training, our method performs only slightly worse than the full vocabulary baseline, which represents what a transformer-based model can do with 10 APs. Our method outperforms both the vanilla and the alpha-renaming baselines by a considerable margin, which is significant since the latter is the only other model that can generalize to more APs. Based on this, we hypothesize that the proposed stochastic AP embeddings provide a more explicit enforcement towards learning embedding-covariant transformations in the model, as opposed to training with alpha-renaming, where the learned embeddings may still carry unwanted token-specific biases. Furthermore, unlike the baseline models, our model does not have to learn the concept of AP from scratch for each AP token thanks to the shared embedding part. This could explain why our method shone against the alpha-renaming baseline in the LTL task where the interchangeable tokens are more complex than the copying task.
Motivation for generalization
@@
The generalization to larger AP counts is important especially when considering the exponential growth of the dataset generation time. In Figure 5, we visualize the growth pattern of the trace checking duration based on increasing formula length and AP count. The times are relative to the fastest trace checking time. The exact times will vary depending on the machine. In our experiments, generating 100000 samples of exact formula length 50 with at most 10 APs took 2 hours and 21 minutes on a system with 56 threads.
Figure 5: Scaling behavior of the trace generation using spot.
Alpha-covariance
@@
We evaluate the alpha-covariance performance of these models in Table 3. Note that since 10 APs lead to a lot more naming permutations than 5 APs, the alpha-covariance values are remarkably smaller compared to Table 2. Unlike the results from Table 2, however, our method outperforms the alpha-renaming approach here. This shows that our method excels in out-of-distribution settings, but trades off some in-distribution performance. Although the full vocabulary baseline performs very similarly to our method, it’s important to note that this region is in-distribution for that model. Overall, these results align with Figure 4.
Alpha-Covariance
Model 3 AP 4 AP 5 AP 6 AP 7 AP 8 AP 9 AP 10 AP
Alpha-Renaming 50.64% 43.00% 40.95% 37.49% 30.80% 30.30% 28.76% 25.57%
Proposed 54.30% 46.05% 45.64% 41.88% 33.89% 35.29% 33.18% 28.34%
@@ 6 Conclusion
The primary difference between machine learning and numerical optimization is the intention to generalize to out-of-distribution samples, for which the network architecture and its inductive biases play a vital role. In this work, we addressed the challenge of generalizing to larger vocabulary sizes unseen during training and creating an inductive bias for alpha-equivalence. We also contributed the alpha-covariance metric for measuring the model’s consistency against alpha-equivalent inputs. These contributions embody a foundation for learning extensible vocabularies for interchangeable tokens, which is especially useful for formal reasoning tasks in which alpha-equivalence naturally arises. Although our dual-part embedding method demonstrates generalization capabilities, its performance in the LTL solving task decreases slightly in in-distribution data (Table 2). The future work can tackle this issue, which may eventually lead to Pareto improvements in bias-variance tradeoff. Moreover, applying our approach to problems in which the interchangeable tokens have meaningful names (e.g., human-written variable names) represents an intriguing area for future research. Finally, new randomization and/or normalization methods for our embeddings can be explored.
@@ Acknowledgments
Appendix A Appendix
A.1 Preliminary: Language models
The autoregressive language modeling or sequence modeling in a broader sense—whose goal is to predict the next token given the past tokens—was revolutionized by the transformer architecture [Vaswani et al., 2017], replacing the step-by-step processing of recurrent neural networks (RNNs) with a parallelizable attention mechanism. At its core lies the attention mechanism, which computes three vectors—query, key, and value—from input embeddings. This mechanism allows the model to weigh the importance of different tokens, enabling it to capture long-range dependencies efficiently. In self-attention, these vectors come from the same sequence, while in cross-attention, key and value vectors come from a different sequence, as in encoder-decoder setups. The transformer consists of an encoder with self-attention and feed-forward layers, and a decoder that adds cross-attention to incorporate the encoder’s output. Since attention lacks an inherent sense of token order, positional encodings are added to input embeddings to provide sequence structure. During training, attention masking ensures causality in predictions, preventing future tokens from being considered when predicting the next one.
A.2 Temporal logic overview
Linear Temporal Logic (LTL) extends propositional logic by introducing the ability to reason about the evolution of propositions over time [Pnueli, 1977]. The syntax of LTL, defined over a finite set of atomic propositions $P$, is as follows: The syntax of LTL, defined over a finite set of atomic propositions $P$, is given in Equation 3, where $\mathbf{T}$ represents True, $p\in P$ an atomic proposition, $\neg$ the negation operator, $\wedge$ the conjunction operator, $\mathbf{X}$ and $\mathbf{U}$ the temporal operators next and until respectively.

### v3

@@ 5 Experiments
Experimental setup. We use a transformer encoder-decoder architecture in all experiments. We always use the same embedding size in both encoder and decoder due to weight tying. We use the RoPE (Su et al., 2024) as the positional encoding method in the decoder. In the encoder, we use tree-positional encoding if applicable (logic tasks), RoPE otherwise (copying task). The hyperparameter settings are given in Table 3 in Appendix D.
Baselines. We train three types of baseline models with traditional embeddings: the first one on the original dataset, the second one on a dataset with the same parameters but using a larger vocabulary size, and the third one on the original dataset but using a data augmentation strategy. Specifically, for the third baseline, the number of interchangeable token embeddings matches that of the test set, and we apply random alpha-renaming at each forward pass during training. This ensures that the model is exposed to all tokens in the test set, but the number of unique interchangeable tokens the model sees in each sample remains limited as in the training set. Note that this is an internal baseline that doesn’t exist in the literature to the best of our knowledge.
@@ 5.1 Copying with Extendable Vocabulary
We introduce a new toy problem designed to evaluate the vocabulary generalization capabilities of our embedding method. We create various training datasets that contain 10 million random strings with a limited vocabulary size. A string is given as input, and the model is expected to produce the input string exactly via autoregressive generation. This embodies a helpful toy problem for our method because all tokens are interchangeable, barring the special tokens (start/end). In these experiments, we expect the model to generalize to larger vocabulary sizes unseen during training.
Using edit distance as our evaluation metric, we first assess the vocabulary generalization capabilities (Appendix E.1). Since our method excels in this task, we then explore generalization in both vocabulary size and string length (Appendix E.2), performing a hyperparameter search over the settings of our embedding method (Appendix E.3). Finally, we scale up the vocabulary size and the string lengths to evaluate our method (Appendix E.5). Our method exhibits perfect performance in the out-of-distribution domain as shown in Figure 2. We also examine our method’s sensitivity to randomness in embeddings (Appendix E.4), and propose using the random embedding with median cross entropy loss as a proxy for average performance.
@@ 5.2 LTL Solving
In this section, we train models on the LTLRandom35 dataset from DeepLTL (Hahn et al., 2021) and other synthetic datasets created with the same method. To evaluate the correctness of the generated formulae, we utilize spot framework version 2.11.6 (Duret-Lutz et al., 2022). We use tree-positional encoding (Shiv & Quirk, 2019) in the encoder and RoPE (Su et al., 2024) in the decoder. We generate predictions using beam search with beam size = 3.
Table 1: Evaluation of the baselines, our method, and Llama 3.2 on the LTLRandom35 dataset. The alpha-renaming baseline was trained using 5 AP embeddings since vocabulary generalization is not evaluated here. First two columns denote the training dataset and the model. Next two columns indicate the ratio of the correct predictions and exact matches on 99,989 test set samples as evaluated by spot. Last three columns display mean alpha-covariance values for varying atomic proposition (AP) counts, evaluated on all alpha-equivalent variants of 1000 test samples. The results indicate that our method induces a robust inductive bias for alpha-equivalence.
@@
Pretrained Llama 3.2 3B 24.33% 0.34% 68.17% 63.27% 62.34%
Baselines. We trained all of the baseline models from scratch. For the first type of baseline, we aimed to reproduce the results from Hahn et al. (2021). Hence, we used the best hyperparameters they reported (Appendix D). Unlike Hahn et al. (2021), we experimented with RoPE (in the decoder) and AdaCos, but did not observe a noteworthy improvement on the validation set. [^1] Using RoPE in the decoder increased the ratio of correct predictions from 97.8% to 98.0% on the validation set. Introducing AdaCos in addition to RoPE increased this value to 98.2%.After determining the best baseline model on the validation set, we evaluated it on the test split of LTLRandom35 and obtained a correct rate of $98.2\%$ against the $98.5\%$ reported by Hahn et al. (2021).
5.2.1 Dataset Perturbations
While the original model performs significantly worse under perturbation, both alpha-renaming and proposed models match the baseline performance in correctness ratio despite perturbation. This observation suggests that these modifications introduce a robust inductive bias that makes the models resistant to perturbations in the data. A minor decrease in the ratio of exact matches is noted, but this may signify less overfitting and a better bias-variance tradeoff in the larger context. Appendix F continues this experiment with limited amount of training samples instead of perturbations.
(a) LTL Solving
(b) Propositional Logic
Figure 3: Heatmaps visualizing the ratio of correct predictions on a special test set, for LTL solving (top) and propositional logic (bottom) tasks. The brightness of the color depends on the sample size, with full brightness representing 100 samples. The dashed white box represents the boundaries of the training dataset. Our model is competitive with the full vocabulary baseline despite being only trained on formulae with at most 5 APs, and outperforms other baselines.
Figure 4: Scaling behavior of the trace generation using spot.
(a) LTL Solving
(b) Propositional Logic
Figure 5: Heatmaps for the ablation studies. The results are reported on the same test set as in Figure 3.
Figure 6: Llama 3.2 heatmaps for the two logic tasks.
Figure 7: Average runtime cost of generating 8-dimensional unique random vectors from Neighboring Points with different uniqueness checking methods.
@@ 5.2.2 Alpha-Covariance
Intuitively, when alpha-covariance is $1$, the model is unaffected by all alpha-conversions in ${\mathbb{P}}$. An alpha-covariance of 00 indicates that $|{\mathbb{U}}|=|{\mathbb{P}}|$, i.e., the model’s prediction for each alpha-equivalent pair is unique after undoing the alpha-conversion. This is unwanted because alpha-conversions should not change the semantic meaning. Thanks to the embedding randomization in our method, an alpha-conversion does not necessarily change the embeddings, and conversely, there are multiple ways to embed the same input.
For the proposed method, we generate the random embeddings once at the start of an evaluation run using the heuristic explained in Appendix E.4. Thus, alpha-conversions in this context are equivalent to shuffling the random embeddings in our method, which amounts to measuring our model’s robustness against differences in random embeddings.
@@
Task Model Alpha-Covariance
3 AP 4 AP 5 AP 6 AP 7 AP 8 AP 9 AP 10 AP
LTL Alpha-Renaming 50.64% 43.00% 40.95% 37.49% 30.80% 30.30% 28.76% 25.57%
Proposed 54.30% 46.05% 45.64% 41.88% 33.89% 35.29% 33.18% 28.34%
Propositional Logic Full Vocabulary 39.77% 30.08% 30.37% 26.64% 20.97% 22.97% 18.80% 17.20%
Alpha-Renaming 42.29% 32.36% 33.45% 30.28% 24.91% 26.47% 22.29% 19.83%
Proposed 43.36% 32.49% 33.65% 30.04% 25.00% 26.63% 21.99% 20.75%
@@ 5.2.3 Generalization
The test dataset for this experiment contains at most 100 formula-trace pairs for each combination of AP count and formula length, whose maximum is 50 instead of 35. We report the results for our model (using Hypercube Vertices, $d_{\beta}=5$) and the three baselines in Figure 3(a). The first baseline uses the same training dataset, whereas the second baseline uses a new LTL dataset with 10 APs, which we create using the same method as LTLRandom35. For the third baseline, we train a fixed embedding model with 10 APs using the same 5 AP dataset but we shuffle the AP embeddings in each forward pass during training.
@@
Discussion. Despite seeing only 5 APs during training, our method performs only slightly worse than the full vocabulary baseline, which represents what a transformer-based model can do with 10 APs. Our method outperforms both the vanilla and the alpha-renaming baselines by a considerable margin, which is significant since the latter is the only other model that can generalize to more APs. Based on this, we hypothesize that the proposed stochastic AP embeddings provide a more explicit enforcement towards learning embedding-covariant transformations in the model, as opposed to training with alpha-renaming, where the learned embeddings may still carry unwanted token-specific biases. Furthermore, unlike the baseline models, our model does not have to learn the concept of AP from scratch for each AP token thanks to the shared embedding part. This could explain why our method shone against the alpha-renaming baseline in the LTL task where the interchangeable tokens are more complex than the copying task.
@@
Motivation for generalization. The generalization to larger AP counts is important especially when considering the exponential growth of the dataset generation time. In Figure 5, we visualize the growth pattern of the trace checking duration based on increasing formula length and AP count. The times are relative to the fastest trace checking time. The exact times will vary depending on the machine. In our experiments, generating 100000 samples of exact formula length 50 with at most 10 APs took 2 hours and 21 minutes on a system with 56 threads.
@@
Alpha-covariance. On the same generalization dataset, we evaluate the alpha-covariance performance of these models in Table 2. Note that since 10 APs lead to a lot more naming permutations than 5 APs, the alpha-covariance values are remarkably smaller compared to Table 1. Unlike the results from Table 1, however, our method outperforms the alpha-renaming approach here. This shows that our method excels in out-of-distribution settings, but trades off some in-distribution performance. Although the full vocabulary baseline performs very similarly to our method, it’s important to note that this region is in-distribution for that model. Overall, these results align with Figure 3(a).
5.3 Assignment Prediction for Propositional Logic
To further demonstrate the applicability and generalization capabilities of our method, we evaluate it on a considerably different logical problem: predicting assignments for propositional logic (Appendix C). The experimental setup is based on DeepLTL (Hahn et al., 2021) with minor differences in hyperparameter choices (Appendix D). We use pyaiger (Vazquez-Chanlatte & Rabe, ) to generate datasets and evaluate predictions. In Appendix G.1, we provide additional details about our experimental setup.
We perform the generalization experiment as in Section 5.2.3 and report the results in Figure 3(b). We observe the same ranking with slightly larger performance gaps. Once more, the proposed method is superior to all approaches that use the same 5 AP training dataset, beaten only by the full vocabulary model which sidesteps the challenge of AP generalization due to its enhanced training dataset.
We continue propositional logic experiments in Table 2 and Appendix G.2, which focus on alpha-covariance and dataset perturbations respectively. The results of these experiments also align with the LTL experiments.
5.4 Ablation Studies
The hyperparameter search in Appendix E.3 operates on the copying task, and, alongside searching over the embedding hyperparameters, experiments with disabling the normalization features and AdaCos, thereby constituting an ablation study. For the LTL and propositional logic tasks, we always kept the normalization features and AdaCos enabled in the previous sections. In this section, we evaluate the impact of these features by disabling them on our best-performing models for these two logic tasks. We ablate one aspect at a time, except for $f_{fn}$, which is disabled together with AdaCos because AdaCos depends on $f_{fn}$ to function correctly.
Figure 5 presents the results, which demonstrate the critical importance of the $f_{bn}$ normalization component. Removing $f_{bn}$ leads to dramatic performance drops (from 90.76% to 29.53% on LTL, and from 77.70% to 14.12% on propositional logic), confirming that maintaining balance between the common and randomized embedding parts is essential for our method’s success. The experiments with AdaCos and $f_{fn}$ indicate task-dependent benefits: they provide significant improvements on LTL (90.76% vs. 81.45% when AdaCos is removed), while showing negligible impact on propositional logic.
5.5 Comparison with LLMs
To contextualize the effectiveness of our proposed approach, we evaluate the performance of a general-purpose LLM (large language model), specifically, the 3B parameter version of Llama 3.2 (Grattafiori et al., 2024), on the LTL task. The details about the prompt design, inference parameters, and implementation are provided in Appendix I.
In the last row of Table 1, we report the performance of Llama 3.2 on the test split of LTLRandom35. These results (e.g., 24.33% correct) are drastically lower than those achieved by our proposed method (95.94%). On propositional logic, Llama 3.2 achieves a slightly better accuracy but much worse alpha-covariance (Table 7 in Appendix G.2). Additionally, we replicate the setups in Figure 3 using Llama 3.2 on the same datasets and sample sizes. As shown in Figure 7, the resulting accuracies are 21.70% (LTL solving) and 30.92% (propositional logic), compared to 90.76% and 77.70% by our method. This striking gap illustrates the limitations of general-purpose LLMs in highly specialized domains such as LTL solving, even when the model size far exceeds that of our dedicated architectures.
5.6 Computational Efficiency
We evaluate the computational efficiency of our method in terms of training time, inference speed, and memory usage (see Appendix H for full details). Our method incurs a modest 13% training overhead compared to the baseline in LTL solving task. At inference, embedding preparation takes only 0.0003 seconds and is required just once at the beginning of an evaluation session, making its cost negligible relative to model execution (0.206 seconds for a forward pass and 9.808 seconds for autoregressive generation). Our optimized method for generating unique random vectors with integer reservoir sampling (Section 4.2) scales efficiently to a large number of vectors unlike the naive approach (Figure 7). While the parameter count of traditional embeddings scales linearly with interchangeable token count, our method’s parameter count remains constant, as embeddings are shared across interchangeable tokens.
6 Limitations
While our method provides an effective framework for enforcing alpha-equivalence in formal languages, it is not directly applicable to natural language, in which tokens carry semantic and contextual information that is often essential for interpretation. For instance, even though variable names like electricity_bill and water_bill may be functionally interchangeable in certain code constructs, they convey distinct meanings that are not preserved under alpha-conversions. As such, enforcing alpha-equivalence may reduce interpretability and degrade performance in tasks that rely on linguistic connotations. This represents an intriguing area for future research.
Another limitation is the requirement to manually define the set of interchangeable tokens, which may not be feasible in some settings. Moreover, our method requires training from scratch due to modifications in the embedding architecture, posing challenges for integration with pretrained models.
Although our dual-part embedding method demonstrates generalization capabilities, its performance in the LTL solving task decreases slightly for in-distribution data (Table 1). The future work can tackle this issue, which may eventually lead to Pareto improvements in bias-variance tradeoff. Finally, new randomization and normalization methods for our embeddings can be explored.
@@ 7 Conclusion
A central goal in machine learning is to generalize to out-of-distribution samples, for which the model design and its inductive biases play a vital role. In this work, we tackle the challenge of generalizing to larger vocabulary sizes unseen during training and creating an inductive bias for alpha-equivalence. We also contribute the alpha-covariance metric for measuring the model consistency against alpha-equivalent inputs. These contributions embody a foundation for learning extensible vocabularies for interchangeable tokens, which is especially useful for formal reasoning tasks in which alpha-equivalence naturally arises.
@@ Acknowledgments
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Preliminary: Language models
The autoregressive language modeling or sequence modeling in a broader sense—whose goal is to predict the next token given the past tokens—was revolutionized by the transformer architecture (Vaswani et al., 2017), replacing the step-by-step processing of recurrent neural networks (RNNs) with a parallelizable attention mechanism. At its core lies the attention mechanism, which computes three vectors—query, key, and value—from input embeddings. This mechanism allows the model to weigh the importance of different tokens, enabling it to capture long-range dependencies efficiently. In self-attention, these vectors come from the same sequence, while in cross-attention, key and value vectors come from a different sequence, as in encoder-decoder setups. The transformer consists of an encoder with self-attention and feed-forward layers, and a decoder that adds cross-attention to incorporate the encoder’s output. Since attention lacks an inherent sense of token order, positional encodings are added to input embeddings to provide sequence structure. During training, attention masking ensures causality in predictions, preventing future tokens from being considered when predicting the next one.
Appendix B Temporal logic overview
Linear Temporal Logic (LTL) extends propositional logic by introducing the ability to reason about the evolution of propositions over time (Pnueli, 1977). The syntax of LTL, defined over a finite set of atomic propositions $P$, is given in Equation 3, where $\mathbf{T}$ represents True, $p\in P$ an atomic proposition, $\neg$ the negation operator, $\wedge$ the conjunction operator, $\mathbf{X}$ and $\mathbf{U}$ the temporal operators next and until respectively.